Determine the electrical force of attraction between two balloons that are charged with the opposite type of charge but the same quantity of charge. The charge on the balloons is 6.6 x 10^-5 C and they are separated by a distance of 150 cm.


1
Expert's answer
2022-03-01T15:18:38-0500

Let's assume, both balloons are perfect spheres. Then the force is given as follows:


"F = k\dfrac{q^2}{r^2}"

where "k = 9\times 10^9N\cdot m^2/C^2, q = 6.6\times 10^{-5}C" and "r = 150cm = 1.5m". Thus, obtain:


"F = \dfrac{9\times10^9\cdot (6.6\times 10^{-5})^2}{1.5^2} \approx 17N"

Answer. 17N.